You are a representative for a cell phone company, and it is your job to promote different cell phone plans.
A. Your boss asks you to visually display three plans and compare them so you can point out the advantages of each plan to your customers.
●Plan A costs a basic fee of Php497.5 per month and Php 1 per text message
● Plan B costs a basic fee of Php4510 per month and has unlimited text messages
● Plan C costs a basic fee of Php2497.5 per month and 5 cents per text message
● All plans offer unlimited calling
● Calling on nights and weekends are free
● Long distance calls are included
B. A customer wants to know how to decide which plan will save her the most money. Determine which plan has the lowest cost given the number of text messages a customer is likely to send.

Answer by Theo(13342)

as far as i can determine, 1 phillipine cent is equal to 1/100 * 1 phillipine peso.

therefore, plan C costs a basic fee of 2497.5 php per month + .05 php per message.

wee now have:

plan A costs 197.5 php per month plus 1 php per text message.
plan B costs 4510 php per month plus 0 php per test message.
plan C costs 2497.5 php per month plus .05 php per text message.

probably the quickest way to figure out which plan is cheaper would be to graph all 3 equations on the same graph.

your graph will look like this.




on the graph, plan A is red, plan B is blue, plan C is green.

from the graph, you can see that:

plan A is cheapest when the number of minutes are less than 24107.737.
plan C is cheapest when the number of minutes are greater than 24107.737.

plan A and plan C cost the same when the number of minutes are 24107.737.

i confirmed by formula that plan A and plan C cost the same at 2104.736842 minutes.
the figure on the graph is rounded.
plan A costs 497.5 + 2104.736842 = 2602.236842
plan C costs 2497 + .05 * 2104.736842 = 2602.236842.
they're the same.

the formula was 497.5 + x = 2497 + .05 * x
subtract .05 * x from both sides of the equation and subtract 497.5 from both sides of the equation to get:
.95 * x = 1999.5
solve for x to get:
x = 1999.5 / .95 = 2104.736842.